OpenStudy (anonymous):

Write an equation in slope-intercept form for the line with slope −2 and y -intercept −5 .

Nnesha (nnesha):

y=mx+B is a slope intercept form where m is the slope and b is y-intercept so just plugin numbers

OpenStudy (anonymous):

y=-2x+-5

Nnesha (nnesha):

looks good but how would you write +-5 with a single sign ?

OpenStudy (anonymous):

y=-2x-5

Nnesha (nnesha):

correct

OpenStudy (anonymous):

thank you so much

Nnesha (nnesha):

y=mx+b is slope intercept form and point slope form is \[\rm y-y_1=m(x-x_1)\]m represent slope and (x_1,y-1) the point line passes through

Nnesha (nnesha):

np